Juniper Jams’ balance sheet shows that its outstanding debt consists of $15,000 in 9 percent notes payable and $48,000 in 6 percent bonds. In addition, Juniper has $60,000 preferred stock that pays a 5 percent dividend. If its marginal tax rate is 40 percent, what is Juniper’s financial breakeven point?
Answer #1

Financial Break Even Point (EBIT) =

Preferred Dividend will be $60,000*5% = $3,000

Interest Expense is on notes payable $15000*9% = $1,350

Interest Expense on Bonds = $48,000*6% = $2,880

Total Interest Expense = $1350+$2880 = $4,230

Financial Break Even point (EBIT) = ($3,000/ (1-.40))+$4,230

=($3,000/0.6)+$4,230

=$5000+$4230

=$9,230

Thus, Juniper's financial breakeven point is $9,230.